The area of a rectangle is 24 square meters. Find the length and width of the rectangle if its length is 5 meters greater than its width. Use an equation and the formula for
Area of a rectangle = (width)(length).

Answer by richwmiller(17219)

L*W=24
L=W+5
W*(W+5)=24
W^2+5w-24=0
(W+8)*(W-3)=0
W=-8 W=3
Measurements must be positive
W=3 L=8